The present age of the father is 42 years and that of his son is 14 years. Find the ratio of
Age of the father to the age of the son, when the son was 12 years old.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The son was 12 years old 2 years ago. So the age father 2 years ago will be
= 42–2 = 40 years
Required ratio = 40 / 12 = (4 × 10) / (4 × 3) = 10 / 3

Thanks